Status Report

NASA Space Station On-Orbit Status 4 Aug 2003

By SpaceRef Editor
August 4, 2003
Filed under , , ,
NASA Space Station On-Orbit Status 4 Aug 2003

All ISS systems continue to function nominally, except as noted previously or below.  Day 100 of Expedition 7’s residency aboard ISS!  (Actually 102 days in space, counting from Soyuz launch).  Underway: Increment 7/Week 14.

Early in the morning, after breakfast (2:40am EDT), FE/SO Ed Lu deployed the acoustic dosimeters for another noise level measuring session (last time done: 7/1-3).  The dosimeters are worn by each crewmember for 24 hours (with a microphone on the shirt collar).  A third dosimeter was deployed in the Lab for a 24-hr. static data take.  [Tonight, after about 12 hours of measurements, data from each dosimeter will be recorded and the hardware power-cycled.  Acoustic data must be taken twice per Increment, each time for the duration of the 16-hour crew work day.]

Afterwards, the Science Officer powered up the MSG (Microgravity Science Glovebox), then supervised another experiment session with the CSLM-2 (Coarsening in Liquid/Solid mixtures #2) experiment.  [In addition to initiating venting of SPU-6 (sample processing unit #6), Ed Lu reprogrammed ECU-2 (electronic control unit #2) in order to change the processing time for SPU-6 to 48 hours.   He also performed a leak check on SPU-6, similar to the recent CSLM-2 vent leak check, except that the water line was included in the decay test.  The leak check should verify that there are no leaks in the water or vent lines as a result of the bent support plate on SPU-6.  If SPU-6 passes the leak check, the sample chamber will continue to be evacuated until 8/6 when Ed will initiate sample processing.]

CDR Yuri Malenchenko prepared equipment assembly and work area for the Russian biomedical “Pilot” experiment (MBI-15) which requires a work table, ankle restraint system and control handles for testing piloting skill.  Yuri then conducted the assessment, followed, for the second time, by flight engineer Ed Lu (last time done: 7/3).  Both were supported by tag-up with ground specialists.  Ed later deactivated, disassembled and stowed the gear.   [Lu performed three flight control modes (fixed, slow and fast free-flyer), each one five times, after checkout and calibration of the control handles.  Results were later reported to the ground.]

Malenchenko performed another 2-hr. inventory audit of 24 Russian medical kits, located behind Service Module (SM) panels 134, 135 and in the medical cabinet niche.  Purpose: to verify their availability, condition and stowage locations.

Lu completed the microbial analysis of the air and surface swab samples collected last week (7/30) with SSK (surface sampler kit) and MAS (microbial air sampler) at the T+5d incubation point.  Microbial specimen data were then loaded into the MEC (medical equipment computer) and the analysis hardware taken down and stowed.  [Because of downmass limitations due to the Shuttle stand-down, the samples cannot be returned to the Microbiology Lab for further analysis on the ground, as usual, but were discarded as wet trash after the analysis.]

Malenchenko worked on the SM water supply system, pressurizing the folded bladders of the Rodnik water storage tanks (BV1; BV2) in the 10P/Progress-247 vehicle, a 3.5 hr. task which is preparatory to the upcoming transfer of liquid waste (urine) to the Rodnik tanks for disposal.       [Each of the two spherical Rodnik tanks consists of a hard shell with a soft membrane (bladder) composed of elastic fluoroplastic, which is used to expel water from the tank by compressed air pumped into the tank volume surrounding the bladder.]

In the Lab and Node, Ed Lu conducted the periodic inspection of the ELPS (emergency lighting power supply) system.  [As a task list item, in the Node Ed had to swap a failed lamp with another one in the Node. Since it took extra time to gain access to this location, the job was allotted 2 hrs. The failed lamp will probably be trashed, and the long-term plan is to replace the current one with one taken from the MPLM the next time it is on orbit.]

Yuri attended to the daily routine maintenance of the SM’s SOZh environment control & life support system as well as the daily preparation of the IMS “delta” file, while Ed performed the regular daily status checkup of autonomously running Lab payloads (PCG-STES010, SAMS, MAMS).

As part of his IMS file preparation task, Malenchenko was asked to locate the KUTZ packing and transport items for the on-board IELKs (individual equipment liner kits, Russian: USIL) for the Soyuz crew return vehicle, last reported to be in the DC-1 docking compartment.  [There are 3 sets, dark grey in color, with a zipper and  lacing, rolled up, labeled with names of Expedition 6 crew members.  Yuri needed to find any 2 KUTZ sets and transfer them to the Soyuz 212/TMA-2 Orbital Compartment (BO).]

Malenchenko also completed the regular inspection of the active BRPK-1 air/liquid condensate separator system in the SM.

Both crewmembers worked out in their daily 2.5-h program of physical exercise, on TVIS treadmill and RED expander, and Ed Lu did the weekly 5-min. routine treadmill maintenance.

Working with experts on the ground, Ed Lu last Friday (8/1) removed the lens filter from the back of the electronic still camera’s 400 mm lens, to expose dust particles that were obscuring photographs taken with the digital camera.  [He was able to clean almost all the dust using airbursts from an empty syringe.  However, the two largest dust particles were inaccessible and could not be cleaned.  The lens would be useful for photographing Shuttle tile during approach for docking.]

The Space Acceleration Measurement System (SAMS) unit has not been recovered yet.  The crew performed an inspection and reported that the unit appears to be powered off.  Additional troubleshooting is expected this week.

On Friday (8/1), MCC-Moscow commanded SM battery #3 to cycling mode.  The battery was then taken offline when parts of the 800A storage unit started to overheat as it cycled.  [The battery reached 42 deg C after 3 hours of cycling operation — the normally expected temperature is 35 deg C.  The normal response for this situation is for TsUP to transition the SM thermal loops (KOV) from KOV-1 to KOV-2 since the latter cools the batteries more efficiently.  But that option wasn’t available since last Thursday’s cooling loop transition problem is still being investigated.] 

Today’s CEO (crew earth observation) optional targets, limited in the current XPOP attitude by flight rule constraints on the use of the Lab nadir/science window, and including the targets of the Lewis & Clark 200-year memorial locations, were Johannesburg, South Africa (opportunity for a panoramic view of the entire Witwatersrand line of cities, right of track.  Numerous small white patches are mine spoil dams and indicate the line of the gold mining towns), Lower Amazon River Basin (mapping swath of those coastlines nearer track was requested), Guangzhou, China (nadir and a touch right. Looking at the head of the long estuary), Western Mediterranean Aerosols (pass just west of Sicily with Etna volcano, which should have been smoking), Lisbon, Portugal (a touch right of track at the head of the largest estuary), and London, England, Great Britain (second mainly cloud-free day.  Looking left of track where the River Thames narrows).
CEO images can be viewed at the websites

See also the website “Space Station Challenge” at

SpaceRef staff editor.